Determine the force in each member of the truss and state if the members are in tension

or compression. Take P = 100 kN.


1
Expert's answer
2021-05-07T08:13:11-0400

Consider the diagram below


"\sum M_B=0 \implies (10*2cos60)+(10*(2+2cos 60))=2R_A \implies R_A =20kN \uparrow"

"\sum F_y=0 \implies R_A+R_B=10+10 \implies R_A+R_B=20 \implies R_B =20-20=0"

Consider point C

"F_{BC}sin60= 10 \implies F_{BC}=11.55kN"

"\sum F_x=0 \implies F_{BC}cos 60=F_{DC} \implies F_{DC}=11.55cos 60=5.78kN"

"\sum F_y=0 \implies F_{AD}sin60=20 \implies F_{AD}=23.09kN"

"\sum F_x=0 \implies F_{AB}=F_{AD}cos60 \implies F_{AB}=11.55kN"